# Reinforcement Learning with Verifiable Rewards: GRPO’s Effective Loss, Dynamics, and Success Amplification

### v2

@@ Abstract.
Group Relative Policy Optimization (GRPO) was introduced in (Shao et al., 2024) and used successfully to train DeepSeek-R1 (Guo et al., 2025) models for promoting reasoning capabilities of LLMs using verifiable or binary rewards. We show in this paper that GRPO with verifiable rewards can be written as a Kullback–Leibler ($\mathsf{KL}$) regularized contrastive loss, where the contrastive samples are synthetic data sampled from the old policy. The optimal GRPO policy $\pi_{n}$ can be expressed explicitly in terms of the binary reward, as well as the first- and second-order statistics of the old policy ($\pi_{n-1}$) and the reference policy $\pi_{0}$. Iterating this scheme, we obtain a sequence of policies $\pi_{n}$ for which we can quantify the probability of success $p_{n}$. We show that the probability of success of the policy satisfies a recurrence that converges to a fixed point of a function that depends on the initial probability of success $p_{0}$ and the regularization parameter $\beta$ of the $\mathsf{KL}$ regularizer. We show that the fixed point $p^{*}$ is guaranteed to be larger than $p_{0}$, thereby demonstrating that GRPO effectively amplifies the probability of success of the policy.
@@ 1. Introduction
Group Relative Policy Optimization (GRPO) was recently introduced in DeepSeekMath (Shao et al., 2024). GRPO closely follows PPO’s optimization framework but differs in how the advantage is computed. Specifically, GRPO estimates the advantage using Monte Carlo rollouts rather than a learned critic. Additionally, GRPO applies whitening to the advantage function, meaning it standardizes the reward’s mean and variance. These statistics are estimated from a “group” of Monte Carlo rollouts corresponding to samples from the LLM policy conditioned on a single input or query to the policy. Whitening the advantage function has been recognized in many PPO implementations as an important ingredient for training stability (Engstrom et al., 2020; Huang et al., 2024). The main novelty in GRPO lies in computing this whitening using Monte Carlo rollouts conditioned on a single input or prompt in the context of LLM training.
@@
Verifiable reward balance simplicity and bias and are thought to be less prone to reward hacking than reward models learned from preference data . Reward hacking is a common issue in reinforcement learning where the policy learns to over-optimize a reward leading to a lower quality of the model (Gao et al., 2023). While verifiable rewards are more resilient to reward hacking, Lambert et al. (2024) showed that for low regularization of the $\mathsf{KL}$ constraint to the reference model, reward hacking occurs when using verifiable constraints. Hence we study in this paper $\mathsf{KL}$-regularized Reinforcement Learning with Verifiable Rewards using GRPO.
@@ 2. GRPO With verifiable Rewards as an Adaptive Weighted Contrastive Loss
where $\mathrm{Var}$ is the variance and for $\epsilon\in[0,1]$
$f_{\epsilon}(x,y)=\min(xy,\text{clip}(x,1-\epsilon,1+\epsilon)y).$
@@ 2.1. GRPO with Clipping
$A(q,o)=\begin{cases}\frac{1-p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=1,\\ -\frac{p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=0.\end{cases}$
which simplifies to :
$A(q,o)=\begin{cases}\sqrt{\frac{1-p}{p}}&\text{if }r(q,o)=1,\\ -\sqrt{\frac{p}{(1-p)}}&\text{if }r(q,o)=0.\end{cases}$
@@
Hence we have conditioning on $q$:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$
$=\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=1}+\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=0}$
$=\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\sqrt{\frac{1-p}{p}}\right)\mathbbm{1}_{r(q,o)=1}+\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},-\sqrt{\frac{p}{(1-p)}}\right)\mathbbm{1}_{r(q,o)=0}$
$=\sqrt{\frac{1-p}{p}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),r(q,o)=1}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)-\sqrt{\frac{p}{(1-p)}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),r(q,o)=0}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)$
@@
Moreover due to the clipping we have:
  • for correct outputs the cost is constant $(1+\epsilon)$ if $\pi_{\theta}(o|q)\geq(1+\epsilon)\pi_{\theta_{\text{old}}}(o|q)$
  • for wrong outputs the cost is $(1-\epsilon)$ if $\pi_{\theta}(o|q)\leq(1-\epsilon)\pi_{\theta_{\text{old}}}(o|q)$,
@@ 2.2. Stabilized GRPO with Clipping
$=\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}\right)\mathbbm{1}_{r(q,o)=1}+\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\frac{-p}{\sqrt{p(1-p)+\varepsilon}}\right)\mathbbm{1}_{r(q,o)=0}$
@@ 2.3. Stabilized GRPO with No Clipping
Note that the formulation (GRPO-No-Clip) is the one currently implemented in HuggingFace TRL library (von Werra et al., 2020), hence we will focus in what follows on this version.
@@
10:         Update the policy model $\pi_{\theta}$ by maximizing the GRPO objective with gradient ascent (Equation (GRPO-No-Clip))
@@ 2.4. GRPO Iterations
and the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in Equation (4). We assume all throughout the paper that
$\pi_{0}=\pi_{\mathrm{ref}}.$
Note that moving the optimization from a parametric space to the probability space can be seen as assuming that the hypothesis class of the parametric policies is large enough to represent all policies.
Note that in GRPO iterations the policy at iteration $n$ depends upon the policy $\pi_{n-1}$ via the probability of success $p_{n-1}$, as well on the reference policy via the $\mathsf{KL}$ regularizer.
@@ 3. GRPO Dynamics: Fixed Point iteration for Probability of Success
  • Monotonicity:
    $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)=-h_{\varepsilon,p_{\mathrm{ref}}}(p)(1-h_{\varepsilon,p_{\mathrm{ref}}}(p))\frac{1-2p}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}$
@@
4. GRPO: Fixed Point Iteration Convergence and Probability of Success Amplification
@@
    Under which conditions on $\beta$ and $p_{\mathrm{ref}}$ do fixed points of GRPO iterations, $p^{*}$, lead to a probability of success $p^{*}$ that is higher than the reference initialization $p_{0}=p_{\mathrm{ref}}$?
@@
Let $10<p_{\mathrm{ref}}<$ 1. Let $p^{*}$ be a fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$ we have $p^{*}>p_{\mathrm{ref}}$, if:
  1. (1)
    $p_{\mathrm{ref}}\leq\frac{1}{2}$ for all $\beta>0$.
  2. (2)
    $p_{\mathrm{ref}}>\frac{1}{2}$ and $\beta\cosh^{2}\left(\frac{1}{2\beta}\frac{1}{\sqrt{\frac{1}{4}+\varepsilon}}\right)\geq\frac{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})(2p_{\mathrm{ref}}-1)}{2[p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon]^{3/2}}.$
Figure 3. Condition for probability amplification on $\beta$ is mostly satisified only for high $p$ and small $\beta$ (Blue area)
We see from Theorem 3 that the fixed point $p^{*}$ of the GRPO iteration leads to an amplification of the probability of success of the reference model $p_{\mathrm{ref}}$. We see from Figure 3 that the condition on $\beta$ is rather mild, and it implies that for high $p_{\mathrm{ref}}$, $\beta$ needs to be selected not too small.
@@
One condition for local convergence is therefore to have:
$|h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p^{*})|=p^{*}(1-p^{*})\frac{|2p^{*}-1|}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}<1$
which is satisfied if :
$\beta>\mathcal{B}(p^{*})=p^{*}(1-p^{*})\frac{|2p^{*}-1|}{2[p^{*}(1-p^{*})+\varepsilon]^{3/2}}.$
@@
Remark 2.
Note that both conditions on $\beta$ are stated conditionally on a prompt $q$, to obtain results uniformly on $q$ we need to take $\sup$ on $q$ in all lower bounds.
@@ 6. Conclusion
Appendix A Proofs of Section 3
@@
The objective in Equation (GRPO Iterations) is concave and hence setting the first order optimality conditions (See for example (Mroueh, 2024) ) we obtain:
@@
Appendix B Proofs of Section 4
@@
This is seen by considering the function
$f(p)=h_{\varepsilon}(p)-p,$
whose zeros correspond to fixed points.
Lemma 2.
For all $\beta,\varepsilon>0$, we have $f(p_{\mathrm{ref}})>0$.
Proof.
We have for all $\beta,\varepsilon>0$ $\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon}}\right)\leq 1$.
$f(p_{\mathrm{ref}})-p_{\mathrm{ref}}$ $=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\,\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon}}\right)}-p_{\mathrm{ref}}$

### v3

@@ Abstract.
Group Relative Policy Optimization (GRPO) was introduced in (Shao et al., 2024) and used successfully to train DeepSeek-R1 (Guo et al., 2025) models for promoting reasoning capabilities of LLMs using verifiable or binary rewards. We show in this paper that GRPO with verifiable rewards can be written as a Kullback–Leibler ($\mathsf{KL}$) regularized contrastive loss, where the contrastive samples are synthetic data sampled from the old policy. The optimal GRPO policy $\pi_{n}$ can be expressed explicitly in terms of the binary reward, as well as the first- and second-order statistics of the old policy ($\pi_{n-1}$) and the reference policy $\pi_{\mathrm{ref}}$. Iterating this scheme, we obtain a sequence of policies $\pi_{n}$ for which we can quantify the probability of success $p_{n}$. We show that the probability of success of the policy satisfies a recurrence that converges to a fixed point of a function that depends on the initial probability of success $p_{\mathrm{ref}}$ and the regularization parameter $\beta$ of the $\mathsf{KL}$ regularizer. We show that the fixed point $p^{*}$ is guaranteed to be larger than $p_{\mathrm{ref}}$, thereby demonstrating that GRPO effectively amplifies the probability of success of the policy.
@@ 1. Introduction
Group Relative Policy Optimization (GRPO) was recently introduced in DeepSeekMath (Shao et al., 2024). GRPO closely follows PPO’s optimization framework but differs in how the advantage is computed. Specifically, GRPO estimates the advantage using Monte Carlo rollouts rather than a learned critic. Additionally, GRPO applies whitening to the advantage function, meaning it standardizes the reward’s mean and variance. These statistics are estimated from a “group” of Monte Carlo rollouts corresponding to samples from the LLM policy conditioned on a single input or query to the policy. Whitening the advantage function has been recognized in many PPO implementations as an important ingredient for training stability (Engstrom et al., 2020; Huang et al., 2024).
@@
Verifiable reward balance simplicity and bias and are thought to be less prone to reward hacking than reward models learned from preference data . Reward hacking is a common issue in reinforcement learning where the policy learns to over-optimize a reward leading to a lower quality of the model (Gao et al., 2023). While verifiable rewards are more resilient to reward hacking, Lambert et al. (2024) showed that for low regularization of the $\mathsf{KL}$ constraint to the reference model, reward hacking occurs when using verifiable constraints. Hence we study in this paper $\mathsf{KL}$-regularized Reinforcement Learning with Verifiable Rewards using GRPO. We note that a recent paper (Vojnovic and Yun, 2025) studies GRPO with a focus on the policy obtained using an approximation of the $\mathsf{KL}$ divergence used in practical implementations.
@@ 2. GRPO With verifiable Rewards as an Adaptive Weighted Contrastive Loss
where $\mathrm{Var}$ is the variance and for $\epsilon\in[0,1]$: $f_{\epsilon}(x,y)=\min(xy,\text{clip}(x,1-\epsilon,1+\epsilon)y).$
@@ 2.1. GRPO with Clipping
$\begin{array}[]{cc}\displaystyle A(q,o)=\begin{cases}\frac{1-p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=1,\\ -\frac{p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=0.\end{cases}&\quad\text{equivalently}\quad\displaystyle A(q,o)=\begin{cases}\sqrt{\frac{1-p}{p}}&\text{if }r(q,o)=1,\\ -\sqrt{\frac{p}{1-p}}&\text{if }r(q,o)=0.\end{cases}\end{array}$
Hence we have conditioning on $q$ and using advantage expressions and $f_{\epsilon}$ for positive and negative advantage we obtain:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$ $=\sqrt{\frac{1-p}{p}}\,\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),\,r(q,o)=1}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)$
$-\sqrt{\frac{p}{1-p}}\,\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),\,r(q,o)=0}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)$
@@ 2.3. Stabilized GRPO with No Clipping
We will focus in what follows on this non-clipped version.
@@
10:         Update the policy model $\pi_{\theta}$ by maximizing (GRPO-No-Clip)) with gradient ascent (Equation
@@ 2.4. GRPO Iterations
and the weights $\omega^{+}_{\varepsilon}$ and $\omega^{-}_{\varepsilon}$ are given in Equation (4). We assume all throughout the paper that $\pi_{0}=\pi_{\mathrm{ref}}.$ Note that moving the optimization from a parametric space to the probability space can be seen as assuming that the hypothesis class of the parametric policies is large enough to represent all policies. Note that in GRPO iterations the policy at iteration $n$ depends upon the policy $\pi_{n-1}$ via the probability of success $p_{n-1}$, as well on the reference policy via the $\mathsf{KL}$ regularizer.
@@ 3. GRPO Dynamics: Fixed Point iteration for Probability of Success
  • Monotonicity: $h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p)=-h_{\varepsilon,p_{\mathrm{ref}}}(p)(1-h_{\varepsilon,p_{\mathrm{ref}}}(p))\frac{1-2p}{2\beta\,[p(1-p)+\varepsilon]^{3/2}}$
@@
4. GRPO: Fixed Point Iteration Convergence and Success Amplification
@@
    Do fixed points of GRPO iterations, $p^{*}$, lead to a probability of success $p^{*}$ that is higher than the reference initialization $p_{0}=p_{\mathrm{ref}}$?
@@
Let $10<p_{\mathrm{ref}}<$ 1. Let $p^{*}$ be a fixed point of $h_{\varepsilon,p_{\mathrm{ref}}}$ we have $p^{*}>p_{\mathrm{ref}}$.
We see from Theorem 3 that the fixed point $p^{*}$ of the GRPO iteration leads to an amplification of the probability of success of the reference model $p_{\mathrm{ref}}$. Note if $p_{\mathrm{ref}}=0$ or $p_{\mathrm{ref}}=1$ , the iteration will lead to $p^{*}=0$ and $p^{*}=1$ respectively.
@@
One condition for local convergence is therefore to have: $|h^{\prime}_{\varepsilon,p_{\mathrm{ref}}}(p^{*})|=p^{*}(1-p^{*})\frac{|2p^{*}-1|}{2\beta\,[p^{*}(1-p^{*})+\varepsilon]^{3/2}}<1$ which is satisfied if : $\beta>\mathcal{B}(p^{*})=p^{*}(1-p^{*})\frac{|2p^{*}-1|}{2[p^{*}(1-p^{*})+\varepsilon]^{3/2}}.$
Remark 2.
Note that the condition on $\beta$ is stated conditionally on a prompt $q$, to obtain results uniformly on $q$ we need to take $\sup$ on $q$ in all lower bounds.
@@ 5. Back to Parametric GRPO Iterations
6. Experimental Validation
Setup
We use the GSM8K dataset from Cobbe et al. (2021) (MIT license), and Qwen2.5-0.5B-Instruct (Apache 2.0 license) by Yang et al. (2024) as the reference policy. We use GRPO implememtation in TRL (von Werra et al., 2020b), and train on the training split of GSM8K on a node with 8 GPUs (GPU0 for the vLLM server and 7 other GPUs for distributed training). We use a learning $5e^{-6}$ and the KL regularizer $\beta=0.1$, and $\mu$ in Algorithm 1 is set to $\mu=10$. Other hyperparameters are given in Appendix B . We use the correctness of the LLM output as a reward.
Success Rate Amplification
The success rate of the policy is then evaluated on the test set consisting of 1319 math questions, where for each question the success rate is evaluated using $50$ samples. We see a success rate amplification from $\pi_{\mathrm{ref}}$ originally (averaged on all prompts) at $21\%$ to $37.5\%$ at the end of the GRPO epoch.
Trajectory of Success rates Along GRPO Iterations
We randomly select few prompts from GSM8K test set and plot in Figure 5(a) the trajectory of the success rate of the model along the GRPO iteration (estimated from 50 samples from the model for each prompt). The success rate is computed from checkpoints of the model along the GRPO training. We see that the trajectory of the success rate $p(q)$ resembles the trajectory of a fixed point algorithm (see Figure 4 in Appendix LABEL:app:plots ). For some points the convegence is fast to the limit point $p^{*}=1$, for others we see an oscillatory behavior (similar to the one in last row in Figure 4). Interestingly when $p_{\mathrm{ref}}=0$, the probability of success does not move much along GRPO iterations as predicted by our theory.
(a) Success rate trajectory of the model on randomly selected prompts along the GRPO iteration.
(b) Convergence to fixed points
Figure 5. Fixed Points like behavior of success rate along GRPO iterations.
Convergence to Fixed Points of $h_{\varepsilon,p_{\mathrm{ref}}}$
We compute the success rate $p_{\mathrm{ref}}(q)$ for all prompts $q$ in the test of GSM8K, and compute the fixed point of $h_{\varepsilon,p_{\mathrm{ref}(q)}}(p)$ for all $q$ using a fixed point algorithm. This gives us $p^{*}(q)$ for all points in the test set. We compute the $\ell_{1}$ error between the success rate along the GRPO training $p_{n}(q)$ for a given prompt and the corresponding fixed point and plot the expected error on the prompts $\mathbb{E}_{q}|p_{n}(q)-p^{*}(q)|$ in Figure 5(b). We see that we have improvement in the convergence to fixed points up to a barrier due to computational and statistical errors as dicussed in Section 5. (Code to reproduce plots is in Supplementary material)
@@ 7. Conclusion
Acknowledgement
We thank Omar El Mansouri for pointing out that no condition on $\beta$ is needed in Theorem 3.
Appendix A Broader Impact and Limitations
Our paper shows the GRPO advantage while simple enjoys a lot of good theoretical properties that ensure success rate amplification when using verifiable rewards. The main limitation in our work is that we study the non-clipped objective. This simplification makes the analysis possible and was used recently in [Vojnovic and Yun, 2025]. Extending the results to clipped objectives or to proximal objectives with $\mathsf{Kl}$ regularization with respect to previous policy remains challenging technically and we hope the techniques introduced in this work will help overcoming those technical challenges. The core contribution being theoretical we don’t see negative societal impact of our work.
Appendix B Assets
Hardware setup
All our experiments were run on one compute node with Dual 48-core Intel Xeon 8468, 2TB of RAM, 8 NVIDIA HGX H100 80GB SMX5, 8x 3.4TB Enterprise NVMe U.2 Gen4, and 10x NVIDIA Mellanox Infiniband Single port NDR adapters, running RedHat Enterprise Linux 9.5.
GRPO Config Setup
We use the group size $G=16$ and per-device batch size $16$ meaning each on each GPU a single prompt $x$ with $16$ corresponding responses is processed. To increase the overall batchsize we use gradient accumulation of $4$, ending with an effective batch size of prompts of $28$. The context length used for this experiment is $200$, and the sampling temperature is set to $\tau=0.1$.
Libraries
Our experiments rely on the open-source libraries pytorch [Paszke et al., 2019] (license: BSD), HuggingFace Transformers [Wolf et al., 2020] (Apache 2.0 license), and HuggingFace TRL [von Werra et al., 2020a] (Apache 2.0 license).
Appendix C Proofs of Section 3
@@
The objective in Equation (GRPO Iterations) is concave and hence setting the first order optimality conditions (See for example [Mroueh, 2024] ) we obtain:
@@
Appendix D Proofs of Section 4
@@
We have for all $\beta,\varepsilon>0$ $\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{p_{\mathrm{ref}}(1-p_{\mathrm{ref}})+\varepsilon}}\right)<1$.
$h_{\varepsilon,p_{\mathrm{ref}}}(p)-p_{\mathrm{ref}}$ $=\frac{1}{1+\frac{1-p_{\mathrm{ref}}}{p_{\mathrm{ref}}}\,\exp\!\left(-\frac{1}{\beta}\frac{1}{\sqrt{p(1-p)+\varepsilon}}\right)}-p_{\mathrm{ref}}$